# {category}

### v6

@@ 2 Bayesian Framework
Whenever possible, we use a relaxed notation leaving implicit the condionalization on the observed dataset, for example, writing $p_{n}(\theta)$ instead of $p_{n}(\theta\,|\,X)$. Further details about the Bayesian framework, including appropriate choices for non-informative reference densities, can be found in [28, 95, 118, 123, 229, 265].
@@ 3 The e-value
In the special case of the flat reference density, $r(\theta)\propto 1$, the surprise function coincides with the posterior probability density function. Its upper cuts, $\overline{T}(v)$, coincide with highest probability density sets (HPDS), making it easier to visualize all the elements defined in the last paragraph. Figure 1 depicts a simple example with a unimodal surprise function and a simply connected tangential set, see [166, Sec.4.3]. However, in general, the posterior probability density function may have multiple local maxima, the tangential set may have multiple connected components, and all these components must be taken into account when computing the e-value. In this respect, the e-value differs from traditional uses of HPDS that only consider a single connected component, usually the connected component containing $\widehat{\theta}$, the unconstrained maximum a posteriory estimator, see [28, 61, 143, 199, 233, 258].
@@ 9 Numerical Example in Model Selection
Figure 7 (left) depicts the polynomial fitting problem for the classical Sakamoto et al. benchmark dataset presented in [198, ch.8]. This dataset, given in Table 1, was generated by a simulation from the i.i.d. (independent and identically distributed) stochastic process
@@
Figure 7: Benchmark (left) and alternative (right) data points, $\circ$, for Sakamoto et al. [198, ch.8] polynomial fitting problem; Exponential target function, $\diamond$; Best fitted polynomials, — , of order 0 to 4; and Order 2 polynomial, *, that renders the smallest regularized error for both datasets.
@@
Table 2: Sakamoto et al. [198, ch.8] benchmark dataset for the polynomial fitting problem
@@
The last columns in Table 3 present, in the context of a linear regression polynomial model of order $k$, the e-value and its standarized version for the statistical hypotheses stating that the parameter of higher order is null, that is, $\,\mbox{ev}\,(H)=\,\mbox{ev}\,(\beta_{k}=0\,|\,y,[x^{0};x^{1};\ldots\,x^{n}])$ and $\mbox{sev}(H)$. This example demonstrates that the e-value and the FBST can replace traditional model selection criteria based on penalized prediction errors. Moreover, using a selection criterion based on the FBST renders a decision process that is invariant, asymptotically consistent, and logically coherent, as explained in the previous sections. In contrast, selection criteria based on penalized prediction errors lack even the most basic invariance properties, even though alternative model selection criteria exist that are invariant and have other desirable statistical properties, see for example [111]. Furthermore, the use of the e-value and the FBST for model selection can be extended to a variety of non-nested (separate) or nested families of hypotheses, including Bayesian classifiers, as analyzed in [19, 35, 106, 135, 136, 137, 138, 157].
@@ 10 Applications
  • Model selection for generalized jump diffusion and Brownian motions, extremal distributions, and persistent memory processes: [6, 18, 134, 183, 184];
@@
  • Testing hypotheses in biological sciences, including cases in ecology, environmental sciences, medical diagnostics, efficacy evaluation of medical procedures, psychology and psychiatry: [2, 35, 127, 142, 153, 169, 200, 204, 217];
@@ 11 Epistemology, Ontology, and Metaphysics
Historically, the development of statistical significance measures (or logical truth-values), and the corresponding tests of hypotheses used in statistical science, have been reciprocally influenced by the epistemological frameworks in which they are presented. Rev. Thomas Bayes (1701-1761), whose work was communicated posthumously by Rev. Richard Price (1723-1791), developed the first methods in this area with clear goals in mind, as stated in [16], also quoted in [11, p.84] and [238, p.245]:
@@
The next two generations in the development of probability and statistics, led by, among others, Pierre-Simon de Laplace (1749-1827) and George Boole (1815-1864), kept these core goals unchanged. While theological questions lost interest over time, the emphasis of statistical research remained essentially metaphysical – in the (gnoseological) sense of searching for and justifying causal explanations for manifested phenomena. Moreover, these causal links were ideally expressed as natural laws in the form of precise or exact quantitative relations, see [229, 230, 238]. Karl Pearson (1857-1936) was the founder and leader of frequentist statistics, the dominant school of thought in statistical science during the XX century. Influenced by the Inverted Spinozism philosophy of Johann Gottlieb Fichte (1762-1814) and the Positivist ideas of Auguste Comte (1798-1857), K.Pearson radically changed the goals of statistical science. He deprecated any form of causal reasoning, the conception or verification of natural laws, or the use of metaphysical (non-observable) entities. Instead, frequentist statistics, following a strict Positivist agenda, only aims to produce good-fitted empirical models able to describe or predict directly observed quantities. Accordingly, the very use of probability calculus is restricted to variables in the sample space, and strictly forbidden for (latent) variables in the parameter space, see [238].
@@
The Objective cognitive constructivism epistemological framework was developed to host the e-value, the FBST, and their formal properties, including their ability to evaluate and test sharp statistical hypotheses. It also provides a naturalized way to ontology and metaphysics in empirical sciences via statistics, that is, a natural way to evaluate empirical support for natural laws and their accompanying causal explanations, and to validate the “objective” use of non-directly observable (or metaphysical) entities, in accordance with the goals of the original works of Bayes and Laplace. For further details on the Objective cognitive constructivism epistemological framework and the way in which, on the one hand, it accommodates the e-value, the FBST and their logical properties, and, on the other hand, it provides a naturalized approach to ontology and metaphysics, see [224, 225, 226, 229, 230, 231, 234, 235, 236, 238, 239, 240, 241, 242].
@@ 12 Future Research and Final Remarks
As already noticed in Section 6, the compositional structure of e-values can be characterized as a possibilistic abstract belief calculus, as analyzed in [27, 50, 51, 60, 63, 27, 131, 222, 233]. This particular algebraic formalism resembles in many ways the formalism used in statistical reliability theory for the analysis of complex systems assembled by serial/ parallel composition of simple(r) elements, see [12, 23, 126, 209]. As noticed in [27, Sec.1], this formal resemblance was a source of inspiration at early stages of this research program. At the same time, the truth-function interpretation of the cumulative surprise function, $W(v)$, was inspired by the work of Ludwig Wittgenstein [48, 262]. Finally, the concept of pragmatic hypotheses and its use in conjunction with the GFBST in testing scientific theories was inspired by sensitivity analysis, as used in optimization and systems’ theory, and its interpretation in terms of fuzzy or paraconsistent logics, see [74, 223]. While the many theoretical results already obtained in this research program justify (in our opinion) the analogies we made and the conceptual links we established between fragments of formal structures used in distinct (and often faraway) research areas, we believe that these intuitive connections deserve and can benefit from an even more rigorous and general setting. In this context, the tools of category theory or other formal abstraction methods, see [36, 37, 84, 141, 218], offer an opportunity for further research in the study of basic logical properties of the aforementioned systems, specially concerning the investigation of their essential compositionality structures and possible generalizations.
@@
In Luhmann’s theory of law, the main purpose of the legal system is congruent generalization of normative behavior expectations, see [149, pp.77,82]. This emblematic statement can be interpreted as follows, see [237]: In Luhmann’s view, norms are neither preexisting conditions nor a priori factual realities. Instead, norms are conceived as intentional projects or idealized models of how society should be; see [149, p.40]. Moreover, in such idealized models, social harmony is based on establishing well-defined, stable, sustainable, and reliable behavioral patterns, also known as eigen-behaviors or eigen-solutions, see [83, 237]. Furthermore, laws and regulations of a society should reflect its norms, stimulating/ penalizing forms of conduct that sustain/ disrupt virtuous eigen-behaviors. Finally, in Luhmann’s view, social norms are not static but essentially dynamic, co-evolving (over long runs) with the behavioral patterns in the society that they simultaneously try to describe, regulate and stabilize (in their present forms, over short runs). In this context, we can try to reinterpret some lessons about model selection and their complexity learned in Section 9, claiming that, as in the case of good empirical laws, good social laws should follow the golden path of equilibrium, avoiding extremes of scarcity and excess, see also [225, 230].
@@
The areas of Logic and Computer science have developed several methods to measure computational complexity, either by counting processing operations in an algorithm or by accessing the code length of its description, see [68, 72, 114, 187, 254, 255]. Some of the methods used to measure complexity in statistical models have already been explained in Section 9. Meanwhile, Information science and Systems’ theory have focused on entropy related measures of complexity, see [1, 10, 64, 123, 229, 232, 243]. In contrast, the mathematical treatment of complexity in theoretical and empirical legal studies is still incipient, see [39, 121, 159, 160, 192, 193, 210, 201, 210]. Naïve adaptations of complexity measures artificially borrowed from other areas often fail to capture relevant aspects of the legal environment.
@@
The asymptotic consistency and standardization procedures presented in Section 5 rely on standard regularity conditions that include continuity and differentiability of several functions defining the statistical model, and also the location of the true value of the parameter at the topological interior of the parameter space. However, many interesting models present non-standard conditions. For example, mixture models and separate hypotheses models often present hypotheses at the (topological) border of the parameter space, see [136, 138], whereas eigenvalue and eigenvector inference problems may present far more challenging conditions, see [202]. The methods presented in [46, 59, 205] and references therein provide tools to generalize the e-value standardization procedure and the corresponding asymptotic results to non-standard regularity conditions.
Other non-standard inference conditions arise in semi-parametric models like, for example, Hilbert space models based on Fourier, orthogonal polynomials, wavelets, and other infinite functional bases. In these cases, it is often necessary to use weakly informative priors that dampen high-frequency modes that effectively reduce the size of resulting truncated models, see [240, Sec.6] and references therein. Further research should explore convenient forms and characteristics of such weakly informative priors and related topics.

### v7

@@
Keywords: e-value, p-value, truth values, statistical significance, Bayesian statistics, hypothesis test, logical composition, epistemology, cognitive constructivism.
@@ 2 Bayesian Framework
Whenever possible, we use a relaxed notation leaving implicit the condionalization on the observed dataset, for example, writing $p_{n}(\theta)$ instead of $p_{n}(\theta\,|\,X)$. Further details about the Bayesian framework, including appropriate choices for non-informative reference densities, can be found in [29, 54, 55, 94, 117, 119, 120, 123, 225, 260].
@@ 3 The e-value
In the special case of the flat reference density, $r(\theta)\propto 1$, the surprise function coincides with the posterior probability density function. Its upper cuts, $\overline{T}(v)$, coincide with highest probability density sets (HPDS), making it easier to visualize all the elements defined in the last paragraph. Figure 1 depicts a simple example with a unimodal surprise function and a simply connected tangential set, see [164, Sec.4.3]. However, in general, the posterior probability density function may have multiple local maxima, the tangential set may have multiple connected components, and all these components must be taken into account when computing the e-value. In this respect, the e-value differs from traditional uses of HPDS that only consider a single connected component, usually the connected component containing $\widehat{\theta}$, the unconstrained maximum a posteriory estimator, see [29, 62, 142, 196, 229, 253].
@@ 9 Numerical Example in Model Selection
Figure 7 (left) depicts the polynomial fitting problem for the classical Sakamoto et al. benchmark dataset presented in [195, ch.8]. This dataset, given in Table 1, was generated by a simulation from the i.i.d. (independent and identically distributed) stochastic process
@@
Figure 7: Benchmark (left) and alternative (right) data points, $\circ$, for Sakamoto et al. [195, ch.8] polynomial fitting problem; Exponential target function, $\diamond$; Best fitted polynomials, — , of order 0 to 4; and Order 2 polynomial, *, that renders the smallest regularized error for both datasets.
@@
Table 2: Sakamoto et al. [195, ch.8] benchmark dataset for the polynomial fitting problem
@@
The last columns in Table 3 present, in the context of a linear regression polynomial model of order $k$, the e-value and its standarized version for the statistical hypotheses stating that the parameter of higher order is null, that is, $\,\mbox{ev}\,(H)=\,\mbox{ev}\,(\beta_{k}=0\,|\,y,[x^{0};x^{1};\ldots\,x^{n}])$ and $\mbox{sev}(H)$. This example demonstrates that the e-value and the FBST can replace traditional model selection criteria based on penalized prediction errors. Moreover, using a selection criterion based on the FBST renders a decision process that is invariant, asymptotically consistent, and logically coherent, as explained in the previous sections. In contrast, selection criteria based on penalized prediction errors lack even the most basic invariance properties, even though alternative model selection criteria exist that are invariant and have other desirable statistical properties, see for example [110, 111]. Furthermore, the use of the e-value and the FBST for model selection can be extended to a variety of non-nested (separate) or nested families of hypotheses, including Bayesian classifiers, as analyzed in [20, 36, 105, 135, 136, 137, 138, 156].
@@ 10 Applications
  • Model selection for generalized jump diffusion and Brownian motions, extremal distributions, and persistent memory processes: [6, 14, 19, 134, 181, 182];
@@
  • Testing hypotheses in biological sciences, including cases in ecology, environmental sciences, medical diagnostics, efficacy evaluation of medical procedures, psychology and psychiatry: [2, 36, 127, 141, 152, 167, 197, 201, 205, 214];
@@
  • Comparison with p-values, other significance measures, and alternative hypothesis tests in miscellaneous applications: [53, 58, 128, 129, 146, 161, 162].
@@ 11 Epistemology, Ontology, and Metaphysics
Historically, the development of statistical significance measures (or logical truth-values), and the corresponding tests of hypotheses used in statistical science, have been reciprocally influenced by the epistemological frameworks in which they are presented. Rev. Thomas Bayes (1701-1761), whose work was communicated posthumously by Rev. Richard Price (1723-1791), developed the first methods in this area with clear goals in mind, as stated in [17], also quoted in [12, p.84] and [234, p.245]:
@@
The next two generations in the development of probability and statistics, led by, among others, Pierre-Simon de Laplace (1749-1827) and George Boole (1815-1864), kept these core goals unchanged. While theological questions lost interest over time, the emphasis of statistical research remained essentially metaphysical – in the (gnoseological) sense of searching for and justifying causal explanations for manifested phenomena. Moreover, these causal links were expressed as natural laws, ideally taking the form of mathematical equations, and later translated as sharp or precise hypotheses in statistical models, see [225, 226, 234]. Karl Pearson (1857-1936) was the founder and leader of frequentist statistics, the dominant school of thought in statistical science during the XX century. Influenced by the Inverted Spinozism philosophy of Johann Gottlieb Fichte (1762-1814) and the Positivist ideas of Auguste Comte (1798-1857), K.Pearson radically changed the goals of statistical science. He deprecated any form of causal reasoning, the conception or verification of natural laws, or the use of metaphysical (non-observable) entities. Instead, frequentist statistics, following a strict Positivist agenda, only aims to produce good-fitted empirical models able to describe or predict directly observed quantities. Accordingly, the very use of probability calculus is restricted to variables in the sample space, and strictly forbidden for (latent) variables in the parameter space, see [234].
@@
The Objective cognitive constructivism epistemological framework was developed to host the e-value, the FBST, and their formal properties, including their ability to evaluate and test sharp statistical hypotheses. It also provides a naturalized way to ontology and metaphysics in empirical sciences via statistics, that is, a natural way to evaluate empirical support for natural laws and their accompanying causal explanations, and to validate the “objective” use of non-directly observable (or metaphysical) entities, in accordance with the goals of the original works of Bayes and Laplace. For further details on the Objective cognitive constructivism epistemological framework and the way in which, on the one hand, it accommodates the e-value, the FBST and their logical properties, and, on the other hand, it provides a naturalized approach to ontology and metaphysics, see [221, 222, 223, 224, 225, 226, 227, 230, 231, 232, 233, 234, 235, 236, 237, 238].
@@ 12 Future Research and Final Remarks
As already noticed in Section 6, the compositional structure of e-values can be characterized as a possibilistic abstract belief calculus, as analyzed in [28, 51, 52, 61, 64, 28, 131, 219, 229]. This particular algebraic formalism resembles in many ways the formalism used in statistical reliability theory for the analysis of complex systems assembled by serial/ parallel composition of simple(r) elements, see [13, 24, 126, 206]. As noticed in [28, Sec.1], this formal resemblance was a source of inspiration at early stages of this research program. At the same time, the truth-function interpretation of the cumulative surprise function, $W(v)$, was inspired by the work of Ludwig Wittgenstein [49, 257]. Finally, the concept of pragmatic hypotheses and its use in conjunction with the GFBST in testing scientific theories was inspired by sensitivity analysis, as used in optimization and systems’ theory, and its interpretation in terms of fuzzy or paraconsistent logics, see [75, 220]. While the many theoretical results already obtained in this research program justify (in our opinion) the analogies we made and the conceptual links we established between fragments of formal structures used in distinct (and often faraway) research areas, we believe that these intuitive connections deserve and can benefit from an even more rigorous and general setting. In this context, the tools of category theory or other formal abstraction methods, see [37, 38, 85, 140, 215], offer an opportunity for further research in the study of basic logical properties of the aforementioned systems, specially concerning the investigation of their essential compositionality structures and possible generalizations.
@@
In Luhmann’s theory of law, the main purpose of the legal system is congruent generalization of normative behavior expectations, see [148, pp.77,82]. This emblematic statement can be interpreted as follows, see [233]: In Luhmann’s view, norms are neither preexisting conditions nor a priori factual realities. Instead, norms are conceived as intentional projects or idealized models of how society should be; see [148, p.40]. Moreover, in such idealized models, social harmony is based on establishing well-defined, stable, sustainable, and reliable behavioral patterns, also known as eigen-behaviors or eigen-solutions, see [84, 233]. Furthermore, laws and regulations of a society should reflect its norms, stimulating/ penalizing forms of conduct that sustain/ disrupt virtuous eigen-behaviors. Finally, in Luhmann’s view, social norms are not static but essentially dynamic, co-evolving (over long runs) with the behavioral patterns in the society that they simultaneously try to describe, regulate and stabilize (in their present forms, over short runs). In this context, we can try to reinterpret some lessons about model selection and their complexity learned in Section 9, claiming that, as in the case of good empirical laws, good social laws should follow the golden path of equilibrium, avoiding extremes of scarcity and excess, see also [222, 226].
@@
The areas of Logic and Computer science have developed several methods to measure computational complexity, either by counting processing operations in an algorithm or by accessing the code length of its description, see [69, 73, 113, 185, 250, 251]. Some of the methods used to measure complexity in statistical models have already been explained in Section 9. Meanwhile, Information science and Systems’ theory have focused on entropy related measures of complexity, see [1, 11, 65, 123, 225, 228, 239, Wallace2005]. In contrast, the mathematical treatment of complexity in theoretical and empirical legal studies is still incipient, see [40, 121, 157, 158, 190, 191, 207, 198, 207]. Naïve adaptations of complexity measures artificially borrowed from other areas often fail to capture relevant aspects of the legal environment.
@@
The asymptotic consistency and standardization procedures presented in Section 5 rely on standard regularity conditions that include continuity and differentiability of several functions defining the statistical model, and also the location of the true value of the parameter at the topological interior of the parameter space. However, many interesting models present non-standard conditions. For example, mixture models and separate hypotheses models often present hypotheses at the (topological) border of the parameter space, see [136, 138], whereas eigenvalue and eigenvector inference problems may present far more challenging conditions, see [199]. The methods presented in [8, 47, 60, 202] and references therein provide tools to generalize the e-value standardization procedure and the corresponding asymptotic results to non-standard regularity conditions.
Other non-standard inference conditions arise in semi-parametric models like, for example, Hilbert space models based on Fourier, orthogonal polynomials, wavelets, and other infinite functional bases. In these cases, it is often necessary to use weakly informative priors that dampen high-frequency modes and that, in this way, effectively shrink resulting truncated models, see [236, Sec.6] and references therein. Further research should explore proper theory and convenient means and methods needed to use the e-value and the FBST under a variety of non-standard regularity conditions.